A cannon fires a round horizontally from the top of a 500 m high mountain with a velocity of 350 m/s.

a. How long (time) does it take the round to hit the ground.

b. How far (distance) from the base of the mountain will it land?
Answer #1

(1) for time taken by round to heat the ground ( eglecting air friction)

h = vy*t - 1/2*g*t^2 vy = vertical speed = 0

-500 = 0*t-1/2*9.81*t^2

t = 10.10 s.

(b) The distance it covers before landing

d = vx*t where vx horizontal speed = 350 m/s

d = 350*10.10

=3533.73 m = 3.53 km.